OpenStudy (anonymous):

what is the solution of the equation 4(y+2)=32.. HELP

OpenStudy (anonymous):

the parenthesis right after the 4 means that you multiply 4 by the whole equation.....meaning you now have 4y + 8 = 32 Then you move the 8 to the other side by subtracting the 8 from both sides. 4y + 8 - 8 = 32 - 8 Simplify and it gives you 4y = 24 Now you divide 4 by both sides: 4y/4 = 24/4 Simplify again and you get y = 6 If this helped please medal and become fan! :)

OpenStudy (anonymous):

exactly what @ChelseaSweets said, distribute and solve for y. Just like she did!

OpenStudy (anonymous):

Thanks :)